# LaserSAM: Zero-Shot Change Detection Using Visual Segmentation of Spinning LiDAR

### v1

@@ IV-A Results
IV-B Limitations
Existing foundation models are trained on colour RGB images, not greyscale nor HSV range-colourized images. The extracted features are likely less expressive than the network’s capability.
@@ IV-B Limitations
Figure 9: Range-coloured image highlighting the effect of the occlusions and submaps. In the red ellipse, a shadow in the scan is visible from the beam of the robot that blocked the LiDAR in the teach. In the magenta ellipse, points are visible in the local blindspot of the sensor because the submap contains a union of points from multiple scans.
@@
Similar to works that process point clouds directly, occluded regions remain a challenge. Figure 9 highlights two regions that are affected by occlusions. Figure 9 is originally rendered from a different pose than the LiDAR meaning that occluded parts of the scene are visible from the new view. This causes a shadow effect from the structure of the robot’s supports. The further from the LiDAR’s origin that the virtual camera is rendered, the more gaps there are in the visual frame. The submap has fewer occlusion-based gaps because it is constructed from sequential frames. Points within the magenta region of Figure 9 were appended to the local submap during the teaching process, filling in some of the holes caused by displacing the virtual camera. If storage was not an issue, retaining every point from nearby teach frames would allow for denser map representations and minimize occlusion effects. In teach and repeat, the robot’s lateral path displacements are typically less than two meters, which minimizes these effects in practice. A final challenge involves the accuracy of the localization algorithm. If the relative transformation between the two point clouds is inaccurate, the rendered images are misaligned and the intersection of the masks check will fail. In the perspective view, small angular misalignments have a large impact on the image location of distant changes. Small translational errors, particularly along the $z$-axis of the camera, impact the position of nearby changes.
IV-C Future Work
@@
Currently, the method has only been evaluated on one type of LiDAR sensor. However, it would be interesting to understand how much the intensity quality impacts overall performance. A potential ablation study could add increasing levels of noise to the intensity data, eventually resulting in uniform white noise coloured only by range. To improve temporal consistency and add tracking, the centroid of previous obstacles will be projected into the camera at the next timestep and used as a prompt. Changes will likely remain visible in the scene, so prompting on their estimated location should improve temporal consistency. Inconsistent masks could be filtered out as spurious. Lastly, this work only considers positive changes: items added in the live scan. It should be possible to detect items that existed in the teach pass on the path that were removed. Due to the manual teach process, it is unlikely that the robot will be driven over items that can disappear so this is left for future investigation.
@@ V Conclusion
Robust environmental perception by mobile robots remains an exciting field for further study. This paper demonstrates that further developments in the hardware sensitivity of LiDAR scanners will enable more algorithms that leverage the benefits of multi-modal detection but use only a single LiDAR sensor. While the resolution of commodity LiDAR-rendered images lags behind conventional cameras, it has reached sufficient thresholds to apply vision algorithms to field robots. It is particularly encouraging that these algorithms naturally work through illumination changes from night to day and day to night. Although the 3.1 Hz Segment Anything Model is slower than the 73 Hz FastSAM, the ability to prompt masks based on the live pixel-level difference in intensity and the previously viewed obstacles makes it more effective than the FastSAM frame-by-frame approach. The classic trade-off between speed and effectiveness leaves both foundation models as viable methods for future evaluation. We look forward to building on these results; implementing the neural network live on the robot and running it in closed-loop with our path planner.

### v2

@@ IV-B Closed-Loop Experiments on Warthog UGV
Figure 8: The change detection quality in the planned corridor between the 3D Difference Baseline and LaserSAM. LaserSAM uses semantic information to capture points closer to the object boundary with the map. This leads to more green pixels on all three objects near the ground.
In the sample trajectory shown in Figure 6, the robot’s obstacle-free path is drawn in blue and is mostly covered by the orange repeat. The robot veers around changes that are detected along the path. Four changes are avoided correctly, and the robot returns to the obstacle-free path afterwards. One small cone was not detected and an emergency stop from the operator was required before continuing. Finally, in one location there was a false positive detection that caused an unnecessary evasive maneuver.
In another experiment, the path was repeated six hours after the mapping was performed after night had fallen. LaserSAM’s performance is unaffected by the ambient illumination of the scene. Figure 7 shows the two virtual images in the afternoon and night. The colouring of the images is qualitatively indistinguishable. This stands in contrast to the conventional camera to which the cone is not visible. An attempt to use SAM for change detection on the RGB camera would fail under these conditions.
The two primary benefits of using segmentation over the baselines are a better definition of changes near the map and the suppression of noisy detections caused by mapping artifacts. Figure 8 shows a pedestrian that is partially segmented based on 3D distances, but using the SAM mask instead, the detection reaches the ground interface. Changes smaller than the 3D baseline’s distance threshold can be detected more accurately with LaserSAM because the intensity data defines their boundaries better than distances alone.
IV-C Limitations
@@
Figure 9: Range-coloured image highlighting the effect of the occlusions and submaps. In the red ellipse, a shadow in the scan is visible from the support strut of the robot that blocked the LiDAR in the teach. In the magenta ellipse, points are visible in the local blindspot of the sensor because the submap contains a union of points from multiple scans.
@@
Similar to works that process point clouds directly, occluded regions remain a challenge. Figure 9 is rendered from a different pose than the LiDAR meaning that occluded parts of the scene are visible from the new view. This causes a shadow effect from the structure of the robot’s sensor supports. The further from the LiDAR’s origin that the virtual camera is rendered, the more gaps there are in the visual frame. In closed-loop testing, this creates an undesirable feedback loop because when the robot deviates from the path to avoid obstacles detection accuracy can decrease. The submap has fewer occlusion-based gaps because it is constructed from sequential frames. Points within the magenta region of Figure 9 were appended to the local submap during the teaching process, filling in some of the holes caused by displacing the virtual camera. If storage was unlimited, retaining every point from nearby teach frames would allow for denser maps and minimize occlusion effects. In teach and repeat, the robot’s lateral path displacements are typically less than two meters, which minimizes these effects in practice.
IV-D Future Work
@@
To improve temporal consistency and add tracking, the centroid of previous obstacles will be projected into the camera at the next timestep and used as a prompt. Changes will likely remain visible in the scene, so prompting on their estimated location should improve temporal consistency. Inconsistent masks could be filtered out as spurious. Lastly, this work only considers positive changes: items added in the live scan. It should be possible to detect items that existed in the teach pass on the path that were removed. In a manual mapping process, it is unlikely that a robot will be driven over items that can disappear so this is left for future investigation.
@@ V Conclusion
This paper demonstrates that developments in the intensity sensitivity of LiDAR scanners enable algorithms that leverage the benefits of multi-modal detection but use only a single LiDAR sensor. While the resolution of commodity LiDAR-rendered images lags behind conventional cameras, it has reached sufficient thresholds to apply vision algorithms to field robots. The zero-shot generalization capabilities of the Segment Anything Model allow for visual change detection to be applied to simulated camera images created from LiDAR data. On the test set, using LaserSAM with 3D-based prompting achieves an IoU of 73.3%. When considering only obstacles within the allowable planning corridor performance is higher with an IoU of 80.4%. With LaserSAM running on a Clearpath Warthog UGV, static changes can be detected and avoided reliably. The operational frequency of 2.8 Hz is fast enough for driving up to 1 m/s. By suppressing spurious false-positive detections the local planner can find better routes. LaserSAM naturally works through illumination changes from night to day and day to night. Continued efficiencies in prompt generation and optimization for inference should help decrease the run time and allow LaserSAM to detect dynamic changes.